Relation between osmotic pressure & relative lowerin of vapour pressure

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Saif
 

Let p0 = Vapour pressure of pure solvent  

p = Vapour pressure of solution  

Pressure at level L1 = p0; Pressure at level L2 = p. Pressure at L1 will be greater than pressure at L2. Then  

P0 – p = hD                        …(ii)  

where ‘d’ is density of solvent vapour of pressure p0.  

Dividing Eq. (i) by (ii), we have,    

\frac{\pi }{p_o -p}=\frac{h\times d}{h\times D}=\frac{h}{D}

at a fixed temperature d/D is constant.  

?    p = constant (p0 - p)             …(iii)  

i.e., p ? (p0 – p)    or\Deltap                …(iv)  

Thus, osmotic pressure is directly proportional to the lowering of vapour pressure.  

 

 

Regards

Arun (askIITians forum expert)